A young man pushes down on a plastic cube with a force of 30.0N. The 8.5 cm high cube is compressed to 5.5 cm. What is the strain on the cube?
1
Expert's answer
2020-05-04T12:41:42-0400
"\epsilon=\frac{\Delta l}{l}=\frac{8.5-5.5}{8.5}=0.353"